OpenStudy (anonymous):

Abdul, while driving to school, computes the average speed for his trip to be 20 km/hr. On his return trip along the same route, there is less traffic and the average speed is 40km/hr. What is the average sped for Abdul's trip?

OpenStudy (kropot72):

Let d = distance from home to school. Time to drive to school = d/20 hours. Time for return trip = d/40 Total time for round trip = d/20 + d/40 = (3*d)/40 Total distance traveled = 2 * d Average speed = \[\frac{2d \times 40}{3d}=\frac{80}{3}km/hr\]

OpenStudy (maheshmeghwal9):

@kropot72 Why my formula isn't used? Please tell:)

OpenStudy (anonymous):

Thankyou thankyou thankyou..!! @maheshmeghwal9 !! ^_^ Thankyou sooo much :*

OpenStudy (maheshmeghwal9):

but thanku for wt? @kropot72 gave the right answer:)

OpenStudy (kropot72):

@maheshmeghwal9 Your method is incorrect. Check my workings and find out the reason.

OpenStudy (maheshmeghwal9):

k:) i gt it @jahanvi I have done wrong.

OpenStudy (maheshmeghwal9):

My formula is use when initial & final velocity is given & acceleration is constant:)

OpenStudy (maheshmeghwal9):

used*

OpenStudy (maheshmeghwal9):

So initial velocity=u final velocity =v So; \[V_{av.}=\frac{u+v}{2}.\]

OpenStudy (maheshmeghwal9):

but in ur question the trip is different So @kropot72 is 100% right:)

OpenStudy (maheshmeghwal9):

@kropot72 gave answer\[\frac{80}{3}km/hr=26.67km/hr.\]which is correct. So don't confuse both fraction & decimal are same^